# Context-aware gate set tomography: Improving the self-consistent characterization of trapped-ion universal gate sets by leveraging non-Markovianity

### v1

@@ III.2 Non-linear amplification, estimation sensitivity and the Fisher information matrix
where $\Sigma(\hat{\bm{\theta}})$ quantifies the precision by the covariance matrix of the estimate [109]. As noted in [94], using the negative log-likelihood $\mathcal{C}_{\rm ML}$ defined in (40), the FI matrix for each of the circuits employed for the long-sequence GST is
@@
To understand the non-linear amplification due to the consecutive phase-space displacements, let us momentarily switch off the SPAM and single-qubit gate errors, and then compute the FI matrix. Motivated by the expected collective dephasing and Ramsey-like experiments [110, 111, 112, 113], we can focus on GST fiducial pairs defined by the initial state $\rho_{0}=\ket{++}\bra{++}$ | and projective measurement on the basis |±=(|$\ket{\pm}=(\ket{0}\pm\ket{1})/\sqrt{2}$. This constitutes a reduced experimental design which, as we show below, contains nonetheless all the information required to enable a prediction of the GST precision. We first compute the spectrum of the Fisher information matrix for this scheme. This can be achieved analytically by means of expression (59), after the calculation of the model probabilities. Then, motivated by the fact that $\mathrm{det}\Sigma(\hat{\bm{\theta}})$ is proportional to the volume enclosed by the covariance elliptical region [75], we use the spectrum of this matrix as an error bound for the two parameters under estimation, in this case $\Gamma_{\rm d}$ and $\Gamma_{\rm th}$. If we plot the square root of each of these eigenvalues as a function of the number of times $p$ the LS gate is repeated, we obtain the error bounds displayed in Fig. 5 with lighter colors. As can be concluded from this figure, the thermal parameter can be estimated with a precision that grows with $1/p^{2}$, which is a substantial gain compared to the $1/p$ gain expected from long-sequence GST. On the other hand, the dephasing parameter ·$\Gamma_{\rm d}$ does not show any context dependence or non-linear sensitivity, and leads to the typical $1/p$ scaling in the estimation precision.
Figure 5: Error bound estimations for GST. We compute the Fisher Information matrix for the LS gate, assuming the initial state $\rho_{0}=\ket{++}\bra{++}$ | and measuring on the same basis, |±$\ket{\pm}=\frac{1}{\sqrt{2}}(\ket{0}+\ket{1})$. This permits us to obtain an estimated scaling for error in both the thermal parameter (blue) and dephasing parameter (red). In lighter colors we plot the estimates resulting from the FI matrix spectrum, while dashed lines represent the estimations obtained from inverting the derivative of the corresponding amplification functions. We modify $\Delta\omega_{L}$ to observe different scaling behaviors, appreciating closing points for the breathing-mode trajectory at different depths. Guides to the aye showcasing different scaling tendencies are plotted in black-dashed lines, demonstrating $1/p$ and $1/p^{2}$ scalings for dephasing and thermal parameter estimates respectively. For the upper panel a, we set $t_{g}\approx 97\mu s$, while for the lower panel b $t_{g}\approx 95\mu s$. At the same time, we adjust $\mathcal{F}_{i,m}^{z}$ to satisfy the maximal-entanglement condition $J_{12}(t_{g})=\pi/4$. Other parameters used are $c=2\times 10^{7}s^{-3}$, $\tau_{c}=5\times 10^{-4}s$ and $\bar{n}_{\mathrm{b}}=5$.
Let us now present a different perspective that allows us to obtain even simpler analytical bounds for the GST precision. For a given point estimation problem in which one aims to estimate a single parameter $\bar{\theta}$, the estimate $\hat{\theta}$ will present an error $\epsilon_{\theta}$. We assume the existence of a function $f(\theta)$ such that the inverse function $f^{-1}\circ f(\theta)=\theta$ exists in a vicinity of the true value $[\bar{\theta}-\delta\theta,\bar{\theta}+\delta\theta]$. Then, if one can estimate $f(\bar{\theta})$ with error $\epsilon_{f}$, an estimate for the target parameter can be obtained by a simple propagation of errors [114], namely
@@ III.3 Context-aware increase of GST precision
Figure 6: Error of parameter estimates as a function of depth for the non-markovian approach. Scaling of the error for both the dephasing parameter (red) and thermal parameter (blue) with depth, computed for the first LS gate execution of the sequence. Guides to the eye with the expected scalings for short depths are displayed for both parameters. The thermal parameter error exhibits $1/p^{2}$ scaling of precision with depth, while respecting the SQL with the number of samples $N_{\mathrm{samples}}$. After the accumulation of free rotations, this tendency is reverted, to the point that no information from the thermal parameter can be obtained when the phase-space trajectory closes. This manifests as divergences of the scaling, given by inverse of the derivative of eq. (63). With the aim of obtaining a reduced variance, for this plot, we sandwich the LS germ with all possible fiducial pairs. In shaded colors, we display the numerical FI error bounds obtained for these fiducial pairs. Parameters used for the simulations are $c=2\times 10^{7}s^{-3}$, $\tau_{c}=5\times 10^{-4}s$ and $\bar{n}_{\mathrm{b}}=5$. Each circuit is sampled $N_{\rm samples}=10^{4}$ times. Light shaded regions represent $95\%$ confidence intervals after averaging over $100$ GST estimations for each depth.
@@
Figure 7: Diamond distance for the light-shift channel as a function of depth for the first LS gate execution. We run GST on simulated data and plot the diamond distance of the estimates as a function of depth. As we are mainly interested in the LS gate scaling, we do not include single-qubit gates in the computation of the diamond distances. We employ an experimental design in which only 12 two-qubit circuits are used in total. We plot the estimations obtained using both the context-dependent version of GST (blue solid markers) and the context-independent one (red solid markers). The $1/p$ scaling is depicted by a black guide to the eye, and is consistent with the results obtained for the context-dependent estimations. On the other hand, context-independent estimations only provide reliable results for short circuits, in which context-dependent effects are still small. Because context-dependence is encoded in the thermal parameter, we also plot the error in the estimation of this parameter with both GST models (empty blue and red markers). For the context-dependent estimation, the thermal parameter error scales as $1/p^{2}$ at short depths. This scaling is also depicted with a black guide to the eye. Parameters used for the simulations are $c=2\times 10^{7}s^{-3}$, $\tau_{c}=5\times 10^{-4}s$ and $\bar{n}_{\mathrm{b}}=5$. Each circuit is sampled $N_{\rm samples}=10^{4}$ times. Light shaded regions represent $95\%$ confidence intervals after averaging over $100$ GST estimations for each depth.
@@ Appendix A Non-Markovianity measure for intermediate maps
$g(t^{\prime})=\lim_{\Delta t\to 0}\frac{1}{\Delta t}\bigg{(}\big{\lVert}\mathcal{E}_{t^{\prime}+\Delta t,t^{\prime}}\otimes\mathbb{1}(\ket{\Phi}\!\bra{\Phi})\big{\rVert}_{1}-1\bigg{)},$ (64)
where, again, |$\ket{\Phi}$ refers to the maximally entangled state on a doubled Hilbert space $\mathcal{H}\otimes\mathcal{H}$, and $\mathcal{E}_{t^{\prime}+\Delta t,t^{\prime}}=\mathcal{E}_{t^{\prime}+\Delta t,0}\mathcal{E}^{-1}_{t^{\prime},0}$ is the instantaneous intermediate map, which might not be completely positive. Typically, we will only be interested in the CP-divisibility of the intermediate LS gates with duration $t_{g}$, and not other partitions of the complete evolution. Thus, it is convenient to define a ‘discrete’ version of the non-Markovianty measure. In analogy to $\mathcal{N}_{\rm CP}$, for a sequence of $p$ LS gates, we define $\mathcal{N}^{d}_{\rm CP}=\sum_{k=1}^{p}g^{d}(k)$, where
$g^{d}(k)=\big{\lVert}\mathcal{E}_{LS}^{(k-1,k)}\otimes\mathbb{1}(\ket{\Phi}\!\bra{\Phi})\big{\rVert}_{1}-1.$ (65)
@@
[…]

### v2

@@ III.2 Non-linear amplification, estimation sensitivity and the Fisher information matrix
[…]
@@
[…]
Let us now present a different perspective that allows us to obtain even simpler analytical bounds for the GST precision. For a given point estimation problem in which one aims to estimate a single parameter $\bar{\theta}$, the estimate $\hat{\theta}$ will present an error $\epsilon_{\theta}$. We assume the existence of a function $f(\theta)$ such that the inverse function $f^{-1}\circ f(\theta)=\theta$ exists in a vicinity of the true value $\bar{\theta}-\delta\theta,\bar{\theta}+\delta\theta]$. Then, if one can estimate $f(\bar{\theta})$ with error $\epsilon_{f}$, an estimate for the target parameter can be obtained by a simple propagation of errors [114], namely
@@ III.3 Context-aware increase of GST precision
Figure 6: Error of parameter estimates as a function of depth for the non-markovian approach. Scaling of the error for both the dephasing parameter (red) and thermal parameter (blue) with depth, computed for the first LS gate execution of the sequence. Guides to the eye with the expected scalings for short depths are displayed for both parameters. The thermal parameter error exhibits $1/p^{2}$ scaling of precision with depth, while respecting the SQL with the number of samples $N_{\mathrm{samples}}$. After the accumulation of free rotations, this tendency is reverted, to the point that no information from the thermal parameter can be obtained when the phase-space trajectory closes. This manifests as divergences of the scaling, given by inverse of the derivative of eq. (63). With the aim of obtaining a reduced variance, for this plot, we sandwich the LS germ with all possible fiducial pairs. In shaded colors, we display the numerical FI error bounds obtained for these fiducial pairs. Parameters used for the simulations are $c=2\times 10^{7}s^{-3}$, $\tau_{c}=5\times 10^{-4}s$ and $\bar{n}_{\mathrm{b}}=5$. Each circuit is sampled $N_{\rm samples}=10^{4}$ times. Light shaded regions represent $95\%$ confidence intervals after averaging over 100100 GST estimations for each depth.
@@
Figure 7: Diamond distance for the light-shift channel as a function of depth for the first LS gate execution. We run GST on simulated data and plot the diamond distance of the estimates as a function of depth. As we are mainly interested in the LS gate scaling, we do not include single-qubit gates in the computation of the diamond distances. We employ an experimental design in which only 12 two-qubit circuits are used in total. We plot the estimations obtained using both the context-dependent version of GST (blue solid markers) and the context-independent one (red solid markers). The $1/p$ scaling is depicted by a black guide to the eye, and is consistent with the results obtained for the context-dependent estimations. On the other hand, context-independent estimations only provide reliable results for short circuits, in which context-dependent effects are still small. Because context-dependence is encoded in the thermal parameter, we also plot the error in the estimation of this parameter with both GST models (empty blue and red markers). For the context-dependent estimation, the thermal parameter error scales as $1/p^{2}$ at short depths. This scaling is also depicted with a black guide to the eye. Parameters used for the simulations are $c=2\times 10^{7}s^{-3}$, $\tau_{c}=5\times 10^{-4}s$ and $\bar{n}_{\mathrm{b}}=5$. Each circuit is sampled $N_{\rm samples}=10^{4}$ times. Light shaded regions represent $95\%$ confidence intervals after averaging over 100100 GST estimations for each depth.
@@ IV Conclusions and outlook
V Code availability
The code to reproduce our findings can be found in GitHub [119].
@@ Appendix A Non-Markovianity measure for intermediate maps
[…]
@@
$D_{m}\coloneq D_{m}\big(\Phi^{\bm{\sigma},\bm{\gamma}}_{m}(t_{g})\textstyle\sum_{l=0}^{p-1}{\mathrm{e}}^{{\mathrm{i}}l\omega_{z,m}t_{g}}+\Phi^{\bm{\sigma},\bm{\gamma}}_{m}(\tau){\mathrm{e}}^{{\mathrm{i}}p\omega_{z,m}t_{g}}\big).$ (66)